### v2

# LongRAG: Enhancing Retrieval-Augmented Generation with Long-context LLMs

Abstract
In traditional RAG framework, the basic retrieval units are normally short. The common retrievers like DPR normally work with 100-word Wikipedia paragraphs. Such a design forces the retriever to search over a large corpus to find the ‘needle’ unit. In contrast, the readers only need to extract answers from the short retrieved units. Such an imbalanced ‘heavy’ retriever and ‘light’ reader design can lead to sub-optimal performance. In order to alleviate the imbalance, we propose a new framework LongRAG, consisting of a ‘long retriever’ and a ‘long reader’. LongRAG processes the entire Wikipedia into 4K-token units, which is 30x longer than before. By increasing the unit size, we significantly reduce the total units from 22M to 600K. This significantly lowers the burden of retriever, which leads to a remarkable retrieval score: answer recall@1=71% on NQ (previously 52%) and answer recall@2=72% (previously 47%) on HotpotQA (full-wiki). Then we feed the top-k retrieved units ($\approx$ 30K tokens) to an existing long-context LLM to perform zero-shot answer extraction. Without requiring any training, LongRAG achieves an EM of 62.7% on NQ and 64.3% on HotpotQA (full-wiki), which is on par with the (fully-trained) SoTA model. Our study offers insights into the future roadmap for combining RAG with long-context LLMs.
LongRAG: Enhancing Retrieval-Augmented Generation
with Long-context LLMs
Ziyan Jiang University of Waterloo ziyanjiang528@gmail.com                        Xueguang Ma University of Waterloo x93ma@uwaterloo.ca                        Wenhu Chen University of Waterloo wenhuchen@uwaterloo.ca
1 Introduction
Figure 1: Traditional RAG vs. LongRAG. (Up) Traditional RAG operates on short retrieval units, where the retriever needs to scan over a massive amount of units to find the relevant piece. In contrast, LongRAG operates on long retrieval units (30x longer). Retriever has a much less workload, which significantly boosts the recall score. LongRAG fully exploits the ability of long-context language models to achieve strong performance.
Retrieval-Augmented Generation (RAG) methods have long been employed to enhance large language models (LLMs) Mialon et al. (2023). Knowledge in the form of natural language can be entirely offloaded from the parametric knowledge of LLMs by leveraging a standalone retrieval component from an external corpus. The existing RAG framework tends to use short retrieval units, such as 100-word passages in popular open-domain question-answering tasks Chen et al. (2017); Lewis et al. (2020); Karpukhin et al. (2020). The retriever is tasked with finding the “needle” (i.e. the precise tiny retrieval unit) from the “haystack” (i.e. the massive corpus with tens of millions of information units). Subsequently, the retrieved units are passed to the reader to generate the final response. On the contrary, the reader only needs to extract answers from these retrievals, which is a fairly easy task. This kind of imbalanced design, with a “heavy” retriever and a “light” reader, puts too much pressure on the retriever. Therefore, existing RAG models Izacard and Grave (2020b) have to recall huge amounts of units, such as the top-100/200, combined with additional re-ranker to achieve the best performance. Moreover, short retrieval units can lead to semantic incompleteness due to document truncation. This can lead to information loss, ultimately hurting the end performance. This design choice was made in an era when NLP models were heavily restricted by their ability to handle long contexts. With the recent advances in long-context language models, the retriever and reader can potentially handle up to 128K or even millions of tokens as input Reid et al. (2024); Achiam et al. (2023).
In this paper, we propose to revisit this design choice for open-domain question answering and propose the LongRAG framework as a solution to balance the workload between the retriever and the reader, as illustrated in Figure 1. There are three important designs in our novel framework:
  1. 1.
    Long Retrieval Unit: By using entire Wikipedia documents or grouping multiple related documents, we can construct long retrieval units with more than 4K tokens. This design could also significantly reduce the corpus size (number of retrieval units in the corpus). Then, the retriever’s task becomes much easier with more complete information.
  2. 2.
    Long Retriever: The long retriever will identify coarse relevant information for the given query by searching through all the long retrieval units in the corpus. Only the top 4 to 8 retrieval units (without re-ranking) are used for the next step.
  3. 3.
    Long Reader: The long reader will further extract answers from the concatenation of retrievals, which is normally around 30K tokens. We simply prompt an existing long-context LM (like Gemini or GPT4) with the question to produce the answers in a zero-shot fashion.
These three novel designs significantly boost the overall performance of RAG on open-domain question-answering tasks like NQ Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). LongRAG has several advantages: 1) It does not require additional re-rankers and the best results can be attained by only considering the top 4-8 retrieved units. 2) The long retrieval unit amalgamates comprehensive information from related documents, which can be used directly to answer multi-hop questions without iterative retrieval.
In our experiments, we adopt off-the-shelf retrievers like BGE Xiao et al. (2023) and readers like Gemini-1.5-Pro (Reid et al., 2024) or GPT-4o OpenAI (2024) without any tuning on NQ or HotpotQA. In our experiments, we reduce the NQ corpus size from 22M to 600K document units, which improves the answer recall@1 from 52% (DPR) to 71%. Similarly, we reduce the HotpotQA corpus size from 5M to 500K, which improves the recall@2 from 47% (DPR) to 72%. The improvement in retriever can significantly benefit the reader model. By exploiting the long-context understanding ability of GPT-4o, LongRAG can achieve an EM of 62% on NQ and 64% on HotpotQA. These results could be comparable to the strongest fully trained RAG models like Atlas Izacard et al. (2022) and MDR Xiong et al. (2020b).
We perform ablation studies in subsection 4.4 to prove why longer retrieval units are necessary. Given a budget of 40K recall tokens, with ‘short retriever units’, we can increase the number of recalled units to reach a marvelously high recall score (91% for recall@200). However, the end performance dips significantly due to the huge amount of ‘hard negatives’, which confuses the reader. With ‘long retriever units’, we observe an entirely different trend. As we recall more units (from 1 to 8 units), both the recall and end performance will increase or plateau. The impact of ‘hard negative’ is much less severe in LongRAG. It shows that LongRAG can better exploit the advances in the long-context LLMs (reader). As the long-context methods evolve, the performance of LongRAG will continue to improve. Therefore, we believe the modern RAG systems should re-consider the granularity of their retrieval units to exploit the advantages of the current long-context LLMs.
Meanwhile, there is still room for improvement in our framework, particularly the need for stronger long embedding models, as shown in Table 3. Additionally, more general methods to formulate long retrieval units beyond hyperlinks will be helpful.
2 Related Work
2.1 Retrieval-Augmented Generation.
Augmenting language models with information retrieved from large corpora has become a popular and effective approach for knowledge-intensive tasks, particularly open-domain question answering. The predominant architecture follows a retriever-reader style Chen et al. (2017); Guu et al. (2020), where the input query retrieves information from a corpus, and a language model uses this information as additional context to make a final prediction. Recent work has focused on improving the retriever (Karpukhin et al., 2020; Xiong et al., 2020a; Qu et al., 2020; Xiong et al., 2020b; Khalifa et al., 2023), enhancing the reader (Izacard and Grave, 2020b; Cheng et al., 2021; Yu et al., 2021; Borgeaud et al., 2022), fine-tuning the retriever and reader jointly (Yu, 2022; Izacard et al., 2022; Singh et al., 2021; Izacard and Grave, 2020a), and integrating the retriever with the black-box language model (Yu et al., 2023; Shi et al., 2023; Trivedi et al., 2022). However, the impact of document granularity on the effectiveness and efficiency of the retrieval-augmented generation pipeline remains underexplored.
2.2 Long Context Large Language Models.
The effectiveness of Transformer-based models is hindered by the quadratic increase in computational cost relative to sequence length, especially when dealing with long context inputs. In order to solve this issue, different approaches have been proposed to mitigate computational issues, including sliding memory window and chunk segmentation (Hao et al., 2022; Ratner et al., 2023; Zhu et al., 2024b). FlashAttention Dao et al. (2022) has also been a pivotal strategy to significantly reduce the memory footprint to almost linear w.r.t sequence length.
To enable length extrapolation, RoPE Su et al. (2021) and AliBI Press et al. (2021) position encodings have shown potential to enable length extrapolation, which have been widely used in the literature. Recent endeavors have explored diverse strategies to tackle this challenge, which is mainly Position reorganization (Jin et al., 2024; An et al., 2024), Position interpolation (Chen et al., 2023a; Peng et al., 2023; Liu et al., 2024). Furthermore, alternative architectures beyond the Transformer have been explored to handle long inputs more naturally. These diverse approaches claim that they can enhance the capabilities of LLMs in processing long context inputs more efficiently.
2.3 Long Context Embedding
Recent efforts also increased the context length for embedding models, extending the supported text snippet length from a limit of 512 tokens to 32k tokens. Typically, the development of long-context embedding models involves first obtaining a long-context backbone model. This can be achieved either by pre-training with long inputs from scratch Günther et al. (2023); Nussbaum et al. (2024); Chen et al. (2024) or by utilizing existing large language models that support longer context Wang et al. (2023). Additionally, some works extend the capabilities of existing embedding models to handle long contexts by applying LLM content window extension methods on embedding models Zhu et al. (2024a); Peng and Quesnelle (2023), or by employing state-space encoder models Saad-Falcon et al. (2024).
3 LongRAG
Our proposed LongRAG framework is comprised of two components: the Long Retriever and the Long Reader. An illustrative example of these two components are depicted in Figure 2.
Figure 2: LongRAG example. On the left side, it shows that the long retrieval unit is grouped by Wikipedia documents through hyperlinks. Each retrieval unit contains an average of 4K tokens, corresponding to multiple related documents. On the right side, it shows a multi-hop question answer test case from HotpotQA. The final result can be achieved by using only a few retrieval units, which is then fed into a long reader.
3.1 Long Retriever
The traditional RAG framework employs smaller retrieval units and prioritizes retrieving the exact fine-grained short context containing the answer. In contrast, our proposed LongRAG framework places greater emphasis on recall, aiming to retrieve relevant context with much coarse granularity. This design choice shifts more burden from the retriever to the reader to extract the exact answers from the relevant context.
We denote our corpus for retrieval as $\mathcal{C}=\{d_{1},d_{2},\ldots,d_{D}\}$, which is a collection of $D$ documents. Formally speaking, the long context retriever is a function: $\mathcal{F}:(q,\mathcal{C})\rightarrow\mathcal{C}_{\mathcal{F}}$ that takes as input a question $q$ and a corpus $\mathcal{C}$ and returns a filtered set of texts $\mathcal{C}_{\mathcal{F}}\subset\mathcal{C}$. In traditional RAG, $\mathcal{C}_{\mathcal{F}}$ is usually small which contains about hundred of tokens, which should contain exact information related to the question $q$. In our framework, $\mathcal{C}_{\mathcal{F}}$ is usually more than 4K tokens, which contains relavant but not exact information related to the question $q$. The long retriever function $\mathcal{F}:(q,\mathcal{C})$ is then divided into three steps:
Formulate long retrieval units
A function is applied to the corpus to form $M$ retrieval units: $\mathcal{G}(\mathcal{C})=\{g_{1},g_{2},\dots,g_{M}\}$. In traditional RAG, the retrieval unit $g$ is typically a short span of passage which is split from the documents $d$, containing hundreds of tokens. In our framework, $g$ could be as long as the whole document or even a group of documents, resulting in much longer retrieval units. We group the documents based on their relationships, using hyperlinks embedded within each document. The grouping algorithm is shown in Algorithm 1. The output group is a list of documents that are related to each other. By having a longer retrieval unit, there are two advantages: First, it ensures the semantic integrity of each retrieval unit; Second, it provides much richer context for tasks that require information from multiple documents.
Algorithm 1 Group Documents Algorithm
Input: $S$ (max group size), $D$ (documents), $\text{adj}[d]$ (related documents for each $d$), $\deg(d)$ (number of related documents for each $d$)
Output: $\mathcal{G}$ (set of groups)
Sort $D$ from low degree ($\deg(d)$) to high degree
Initialize an empty set of groups $\mathcal{G}$
for each document $d$ in $D$ do
[…]
     for each related document $r$ in $\text{adj}[d]$ do
         for each group $g$ in $\mathcal{G}$ do
              if $r\in g$ then
[…]
              end if
         end for
     end for
     Create a new group $g_{\text{new}}=\{d\}$
     Sort related_groups by their size
     for each group $g$ in related_groups do
         if $|g_{\text{new}}|+|g|\leq S$ then
              $g_{\text{new}}\leftarrow g_{\text{new}}\cup g$
              Remove $g$ from $\mathcal{G}$
         end if
     end for
     Add $g_{\text{new}}$ to $\mathcal{G}$
end for
return $\mathcal{G}$
Similarity search
We utilize an encoder, denoted as $E_{Q}(\cdot)$, to map the input question to a $d$-dimensional vector. Additionally, we employ a different encoder, $E_{C}(\cdot)$, to map the retrieval unit to a $d$-dimensional vector. We define the similarity between the question and the retrieval unit using the dot product of their vectors:
$sim(q,g)=E_{Q}(q)^{T}E_{C}(g)$
In LongRAG settings, $E_{C}(g)$ is challenging given the length of $g$, so we resort to an approximation as below.
$sim(q,g)$ $=E_{Q}(q)^{T}E_{C}(g)$
$\approx\max_{g^{\prime}\subseteq g}(E_{Q}(q)^{T}E_{C}(g^{\prime}))$
We approximate it by maximizing the scores of all chunks $g^{\prime}$ within the retrieval unit $g$, akin to the MaxP design in Dai and Callan (2019). We consider different levels of granularity, including passage level, document level, and the complete grouped document. The empirical study about this settings is in Table 3. With this similarity score setup, we will retrieve the top $k$ retrieval units closest to the given query. For efficient retrieval, we precompute the embedding of each retrieval unit $g^{\prime}$ and predict the exact inner product search index in FAISS (Johnson et al., 2019).
Aggregate retrieval result
We will concatenate the top $k$ retrieval units into the long context as the retrieval result, denoted by $\mathcal{C}_{\mathcal{F}}=\text{Concat}(g^{1},g^{2},\ldots,g^{k})$. Depending on the selection of retrieval units, a larger retrieval unit size will result in a smaller value of $k$ being used. For instance, if the retrieval unit is a passage, $k$ is approximately above 100; if it’s a document, $k$ is around 10; and for grouped documents as retrieval units, we typically set $k$ to 4 to 8.
3.2 Long Reader
The long reader operates straightforwardly. We feed the related instruction $i$, the question $q$, and the long retrieval result $\mathcal{C}_{\mathcal{F}}$ into an LLM, enabling it to reason over the long context and generate the final output. It’s important that the LLM used in the long reader can handle long contexts and does not exhibit excessive position bias. We select Gemini-1.5-Pro Reid et al. (2024) and GPT-4o OpenAI (2024) as our long reader given their strong ability to handle long context input.
We utilize different approaches for short and long contexts. For short contexts, typically containing fewer than 1K tokens, we instruct the reader to directly extract the answer from the provided context retrieved from the corpus. For long contexts, typically longer than 4K tokens, we empirically find that using a similar prompt as for short contexts, where the model extracts the final answer directly from the long context, often leads to decreased performance. Instead, the most effective approach is to utilize the LLM as a chat model. Initially, it outputs a long answer, typically spanning a few words to a few sentences. Subsequently, we prompt it to generate a short answer by further extracting it from the long answer. The prompt is provided in the Appendix 6.1.
4 Experiments
In this section, we will first detail the dataset we adopt, and then demonstrate the retriever performance. Finally, we will show the end question-answering performance.
Retrieval Unit Corpus Size Num of Retrieval Units Average Num of Tokens Answer Recall (AR)
Corpus Test Set
Passage 22M 1 120 130 52.24
100 12K 14K 89.92
200 24K 28K 91.30
Document 3M 1 820 4K 69.45
5 4K 18K 85.37
10 8K 34K 88.12
Grouped Documents 600K 1 4K 6K 71.69
4 16K 25K 86.30
8 32K 50K 88.53
Table 1: The table illustrates the retrieval performance on NQ. Employing a long-context retriever (with an average number of tokens for each retrieval unit up to 6K) compresses the corpus size by up to 30 times (from 22M to 600K), enhancing top-1 answer recall by approximately 20 points (from 52.24 to 71.69). Furthermore, long-context retrieval requires significantly fewer retrieval units (10 times fewer) to achieve comparable results. Therefore, integrating long-context retrieval significantly alleviates the burden on the retriever model.
4.1 Data
Our proposed methods are tested on two Wikipedia-related question answering datasets: Natural Questions and HotpotQA.
Natural Question
(Kwiatkowski et al., 2019) was designed for end-to-end question answering. The questions were mined from real Google search queries and the answers were spans in Wikipedia articles identified by annotators. This dataset contains 3,610 questions.
HotpotQA
(Yang et al., 2018) consists of two-hop questions over diverse topics. We focus on the fullwiki setting in which two Wikipedia passages are required to answer the questions. Since the gold passages for the test set are not available, we follow prior work (Xiong et al., 2020b) and evaluate on the development set, which has 7,405 questions. There are two main question types in HotpotQA: (1) comparison questions usually require contrasting two entities and (2) bridge questions can be answered by following a connecting entity that links one document to another.
Wikipedia (Knowledge Source)
We use different versions of English Wikipedia for different datasets following previous works Lewis et al. (2020); Yang et al. (2018). For NQ, we use the Wikipedia dumps from December 20, 2018, which contain approximately 3 million documents and 22 million passages. For HotpotQA, we use the abstract paragraphs from the October 1, 2017 dump, which contain around 5 million documents. For each page, only the plain text is extracted and all structured data sections such as lists, tables and figures are stripped from the document.
4.2 Retrieval Performance
Metrics
Retrieval performance is measured using Answer Recall (AR) and Recall (R). For NQ, we use only answer recall, while for HotpotQA, we use both metrics. Answer Recall is the recall of the answer string in all the retrieved documents that we plan to use in the reader. For example, if the retrieval unit is at the “passage” level and the number of retrieval units is 100, answer recall measures whether the answer string is present in these 100 passages. For HotpotQA, we compute AR only for questions with span answers, specifically the “bridge” type questions, while ignoring yes/no and comparison questions, following previous work (Khalifa et al., 2022). Recall used for HotpotQA measures whether the two gold documents are present in all the retrieved results. For example, if the retrieval unit is at the “document” level and the number of retrieval units is 10, recall measures whether both gold documents are present among the 10 retrieved documents.
Experiment Setup
We leverage open-sourced dense retrieval toolkit, Tevatron Gao et al. (2022), for all our retrieval experiments. The base embedding model we used is bge-large-en-v1.5, a general-purpose embeddings model that isn’t specifically trained on our test data.
Retrieval Unit Corpus Size Num of Retrieval Units Average Num of Tokens Recall (R) Answer Recall (AR)
Corpus Test Set
Document 5.2M 2 130 200 30.01 47.75
100 6.5K 10K 74.84 84.67
200 13K 20K 79.68 88.34
Grouped Documents 500K 2 1K 8K 56.30 72.49
8 4K 29K 74.71 84.40
Table 2: The table illustrates the retrieval performance on HotpotQA. Similar to the findings on NQ, a long-context retrieval could significantly alleviate the burden on the retriever component within the entire RAG framework.
Table 1 and Table 2 have shown the retrieval results on NQ and HotpotQA. In the NQ dataset, we utilize three different retrieval units, ranging from shorter to longer: passage, document, and grouped documents. In the table, we have mentioned two kinds of average number of tokens in each retrieval unit: one for the entire corpus and one for each test set. The retrieval units for each test case can sometimes be much longer than the average size across the whole corpus, as the corpus might include some Wikipedia pages with very few words, while the test cases may focus more on longer documents. Generally, our long-context retriever (at the document level and grouped document level) uses retrieval units containing an average of 6K tokens. By using longer retrieval units, there are several advantages: 1) It will significantly alleviate the burden on the retriever by compressing the corpus size by approximately 30 times, from 22M to 600K. The top-1 answer recall improves by about 20 points, from 52.24 to 71.69. We could use significantly fewer retrieval units to achieve comparable retrieval performance. For instance, 8 retrieval units at the grouped document level can achieve similar recall as 100 retrieval units at the passage level. 2) It could provide more comprehensive information to the reader. In the original passage-level RAG setup, information might be incomplete due to the chunking operation. In the HotpotQA dataset, we observe similar results. One notable difference is that in HotpotQA, the retrieval units are only at the document level and grouped document level, as HotpotQA uses only abstract paragraphs from each Wikipedia page.
Model Granularity AR@1
BGE-Large 512-tokens chunk 71.7%
E5-Mistral-7B 4000-tokens chunk 54.2%
E5-Mistral-7B entire grouped retrieval unit 23.4%
Table 3: Different methods to encode the long retrieval unit in the long retriever. Using a general embedding model and approximating by maximizing the similarity scores between the query and all chunks within the retrieval unit is better than using the long embedding model to encode the entire context.
Encode the long retrieval unit
As discussed in Section 3.2, it’s very challenging to employ an encoder, $E_{C}(\cdot)$, to map the retrieval unit $g$ to a $d$-dimensional vector when $g$ is very long. Therefore, we use an approximation in our proposed system. Table 3 demonstrates that our approximation, $sim(q,g)=E_{Q}(q)^{T}E_{C}(g)\approx\max_{g^{\prime}\subseteq g}(E_{Q}(q)^{T}E_{C}(g^{\prime}))$, is much more effective than encoding the entire long context directly. We compare three methods: 1) Using the general embedding model “bge-large-en-v1.5” (Xiao et al., 2023), with $g^{\prime}$ selected as text of 512-token size. 2) Using long embedding model “E5-Mistral-7B” (Zhu et al., 2024a), with $g^{\prime}$ selected as the whole document, which has an average size of 4K tokens. 3) Using long embeddings model “E5-Mistral-7B”, with no approximation, encoding the entire $g$ directly, where $g$ has an average size of 6K tokens. We can notice from the table that our approximation by taking the maximum score between the query and each text piece from the long context produces much better results than encoding them directly using the long embedding model. We believe future improvements in the research direction of long embedding models will further enhance our framework to reduce memory consumption.
Method EM
Closed-Book
GPT-4-Turbo Achiam et al. (2023) 41.2
Gemini-1.5-Pro Reid et al. (2024) 47.8
Claude-3-Opus Anthropic (2024) 49.2
Fully-supervised RAG
REALM Guu et al. (2020) 40.4
DPR Karpukhin et al. (2020) 41.5
RAG Lewis et al. (2020) 44.5
RETRO Borgeaud et al. (2022) 45.5
RePAQ Lewis et al. (2021) 47.8
Fusion-in-Decoder (Izacard and Grave, 2020b) 51.4
EMDR2 Singh et al. (2021) 52.5
Atlas (Izacard et al., 2022) 64.0
No Fine-tuning RAG
REPLUG (Shi et al., 2023) 45.5
LongRAG (Gemini-1.5-Pro; Recall 4 units) 58.6
LongRAG (GPT-4o; Recall 4 units) 62.7
Table 4: The table shows the QA results on the NQ dataset. We compare the results with three groups of baselines: closed-book, which involves directly prompting state-of-the-art LLMs with 16-shot in-context examples; fully-supervised RAG, where the RAG framework is used and the model is fully supervised and trained on the training data; and No Fine-tuning RAG, which employs the RAG framework without any tuning.
Method EM
Closed-Book
Claude-3-Opus Anthropic (2024) 32.8
Gemini-1.5-Pro Reid et al. (2024) 33.9
GPT-4-Turbo Achiam et al. (2023) 42.4
Fully-supervised RAG
CogQA Ding et al. (2019) 37.1
DrKIT Dhingra et al. (2020) 42.1
Transformer-XH Zhao et al. (2019) 51.6
QAMAT+ Chen et al. (2023b) 57.6
HGN Fang et al. (2019) 59.7
PathRetriever Asai et al. (2019) 60.0
HopRetrieve Li et al. (2021) 62.1
MDR Xiong et al. (2020b) 62.3
HopRetrieve-plus Li et al. (2021) 66.5
AISO Zhu et al. (2021) 68.1
COS Ma et al. (2023) 68.2
No Fine-tuning RAG
DSP Khattab et al. (2022) 51.4
PromptRank Khalifa et al. (2023) 55.7
LongRAG (Gemini-1.5-Pro; Recall 8 units) 57.5
LongRAG (GPT-4o; Recall 8 units) 64.3
Table 5: The table shows the QA results on the Hotpot-QA dev set. We compare the results with three groups of baselines: closed-book, which involves directly prompting state-of-the-art LLMs with 16-shot in-context examples; fully-supervised RAG, where the RAG framework is used and the model is fully supervised and trained on the training data; and No Fine-tuning RAG, which employs the RAG framework without any tuning.
4.3 Full QA Performance
We leverage Gemini-1.5-Pro and GPT-4o as the reader in our LongRAG framework. The prompt we use for our experiments are in Table 6. We also refine the standard exact match rate definition to more fairly evaluate LongRAG’s performance. More details can be found in Section 6.2.
We compare our model with several groups of strong previous models as baselines. The first group is “Closed-Book”: These baselines mean that no retrieval component is used; instead, state-of-the-art LLMs are employed to directly obtain the final result. We evaluate our results on Gemini-1.5-pro Reid et al. (2024), Claude-3-Opus Anthropic (2024) and GPT-4-Turbo Achiam et al. (2023). All models are evaluated on 16-shot in-context learning with direct prompting; The second group is “Fully-supervised RAG”, and these baselines involve full-supervised fine-tuning on the training dataset. The third group is “No Fine-tuning RAG”, and these baselines doesn’t involve any supervised fine-tuning on the training dataset.
The QA results on NQ are presented in Table 4, and the QA results on HotpotQA are presented in Table 5. On the NQ dataset, LongRAG achieves a 62.7 exact match rate, which is on par of the strongest fine-tuned RAG model like Atlas. On the HotpotQA dataset, LongRAG achieves a 64.3 exact match rate, which is also close to the SoTA fully-supervised RAG frameworks.
Figure 3: This figure compares different settings of LongRAG on the NQ dataset. This table leverages 200 test cases from the test set to help compare different retrieval unit selections and optimal number of retrieval units fed into the reader (Gemini-based).
Figure 4: This table compares different settings of LongRAG on the HotpotQA dataset. This table leverages 200 test cases from the test set to help compare different retrieval unit selections and the optimal number of retrieval units fed into the reader (Gemini-based).
Figure 5: This table shows the end performance of different granularity of retrieval units. The green curve indicates the ‘large unit’ adopted in LongRAG. The number above the curve indicates the recall. It can be seen that end-performance does not increase monotonically with recall score.
4.4 Ablation Studies
We perform several in-depth ablation to understand what are the important factors in our LongRAG system including "unit size" and "reader variant".
Retrieval Unit Selection
Figure 3 and Figure 4 compare different settings of LongRAG. This table leverages 200 random test cases from the test set to help compare different retrieval unit granularity selection and the optimal number of retrieval units used in the reader. On the NQ dataset, we have two observations: First, regardless of which retrieval unit is selected, there will be a turning point where feeding more retrieval units into the reader becomes detrimental. This is due to the excessive burden placed on the reader, preventing it from effectively understanding and extracting relevant information from the long context. For passage-level retrieval units, the turning point is between 100 and 200; for document-level retrieval units, the turning point is between 5 and 10; and for grouped documents level, the turning point is between 4 and 8. In general, the most suitable context length fed into the reader is around 30K tokens. Second, the semantic integrity is important when comparing the performance of passage-level retrieval units with document or grouped documents level retrieval units, highlighting the advantage of using longer and more complete retrieval units.
Recall vs. EM
In Figure 5, we compare the relationship between retrieval recall and end performance across varying context lengths for different retrieval unit selections. We observe that using fewer retrieval units in the reader with longer retrieval units design reduces the introduction of distractors or hard negatives under a given length budget. Consequently, the end performance does not increase monotonically with the recall score. In the future, with advancements in long embedding models and improved retrieval recall for long retrieval units, we can expect better end performance.
Figure 6: This figure compares different readers of LongRAG on the NQ dataset. This table leverages 200 test cases from the test set to help compare performance using different readers.
Reader Model
In Figure 6, we compare the performance of six different readers: Gemini-1.5-pro, GPT-4-Turbo, GPT-4o, Claude-3-Opus, Claude-3.5-Sonnet and DeepSeek-V2-Chat. The results indicate that GPT-4o achieves the highest exact match score on the 200 test questions of the NQ dataset among the three models. This suggests that GPT-4o is the most effective in the role of a long reader in the LongRAG framework. The enhanced performance of GPT-4o can be attributed to its superior ability to process and comprehend lengthy contexts, ensuring that crucial information is accurately extracted. Therefore, we mainly report the GPT-4o results in our main table. Besides, Gemini-1.5-pro, GPT-4-Turbo, Claude-3-Opus, and Claude-3.5-Sonnet could achieve very similar results. These state-of-the-art black box LLMs are also effective readers within the LongRAG framework. Deepseek-V2-Chat is one of the best open-source LLMs, but its performance degrades significantly compared to the previous five black-box LLMs. The above experiments demonstrate that our current framework depends on the long-context understanding ability of LLMs, and we still have a long way to go in harnessing open-source LLMs within our framework.
5 Conclusion
In this paper, we propose a new framework, LongRAG, to alleviate the imbalance between the burden of the retriever. The LongRAG framework consists of a “long retriever” and a “long reader” component on top of the 4K-token retrieval units. Our proposed framework can significantly reduce the corpus size by 10 to 30 times, which greatly improves the recall of the retriever. On the other hand, the long retrieval unit preserves the semantic integrity of each document. We test our framework on end-to-end question answering tasks and demonstrate its superior performance without any training. We believe LongRAG can pave the road for the modern RAG system design.
Limitation
There are three major limitations of our proposed framework. First, it relies on the long embedding model. Although recent studies have made progress in this direction, there is still a need for stronger long embedding models. In our work, we use an approximation to calculate the semantic score with a regular embedding model, which proves more effective than using a long embedding model. Future improvements in long embedding models could help us further enhance the performance of our system and reduce the storage size of corpus embeddings if the entire long context could be encoded directly. The second limitation is that we only use a black-box LLM as the reader. A reader that supports long input and is less affected by position bias is necessary. Currently, most open-source LLMs do not meet these requirements. The third limitation is that our grouping methods are based on hyperlinks, which are specific to the Wikipedia corpus. A more general grouping method should be considered.